# WaveStitch: Flexible and Fast Conditional Time Series Generation with Diffusion Models

### v2

@@ 5. Evaluation
  • RQ3: Does the conditional loss improve generation quality compared to using only self-guidance or RePainting (Lugmayr et al., 2022; Chen et al., 2023)?
  • RQ4: What is the impact of alternative conditional loss formulations on the generation quality?
  • RQ5: How well does WaveStitch generalise to random imputation tasks with varying missingness ratios?
System Configuration. We use a 12-core AMD Ryzen 9 5900 processor with an NVIDIA RTX 3090 GPU. For software, the CUDA version is 12.1, Ubuntu OS version 20.04.6, and Pytorch version 2.2.1 with Python 3.12.
@@ 5.1. Datasets and Evaluation Framework
Datasets. We use public multivariate and univariate datasets from diverse domains: BeijingAirQuality (BQ) (Chen, 2017), MetroTraffic Volume (MT) (Hogue, 2019), PanamaEnergy (PE) (Aguilar Madrid, 2021), RossmanSales (RS) (FlorianKnauer and Cukierski, 2015), and AustraliaTourism (AT) (Blanche, 2020). Each dataset includes metadata, grouped in the order specified in Table 2, reflecting how the data is organised. The time series channels for each dataset are as follows: for BQ, we consider all gas sensor readings, temperature, pressure, dewpoint, rainfall, and wind speed from the first six stations, sorted alphabetically; for MT, we use temperature, rainfall, snowfall, traffic volume, and cloud cover levels; for PE, we include energy data such as T2M, QV2M, TQL, and W2M; for RS, we focus on sales and customer totals from stores 1-10; and for AT we use the number of trips.
Evaluation Tasks. Creating train-test splits for our problem scenario is challenging. A random split would mix contexts, disrupting dependencies and evaluation integrity. Traditional splits like 80-20 also risk fragmenting temporal dependencies. To address this, we place all samples with a shared root-level value (e.g., a specific year) in the test set, ensuring purely unseen metadata combinations. This approach supports coarse-grained tasks and also allows imposing fine-grained or mid-level tasks by specifying lower-level metadata features within the test set.
To evaluate the model’s performance across different scenarios, we use a structured approach that conditions metadata at three levels: Root (R), Intermediate (I), and Bottom (B), each corresponding to a different level in the grouping hierarchy and reflecting varying degrees of missing signal information and task granularity (see Table 2). Root Level (R) conditions fix the root-level metadata feature (e.g. Year in a Year-¿Month-¿Day hierarchy), creating tasks missing contiguous signal blocks (e.g., 4(a)). Intermediate Level (I) tasks restrict intermediate-level metadata features, giving missing blocks of intermediate size (e.g., 4(b)). Finally, Bottom Level (B) tasks restrict the last metadata feature, yielding fine-grained imputation scenarios (e.g., 4(c)). As shown in Table 2, both I and B additionally condition on the root level feature to ensure samples remain within the test set. While these configurations represent only a subset of possible scenarios, they cover a broad range of missingness patterns, allowing us to evaluate the model under various conditions. For example, this setup can also capture random missingness by combining multiple (B) conditions.
@@ 5.2. Model Implementation and Baselines
We extend the implementations of several baselines from Table 10 for conditional generation. For all diffusion models, including ours, we use SSSD (Alcaraz and Strodthoff, 2023) as the denoiser backbone for consistency, which utilises structured state space layers (S4 layers (Smith et al., 2022)) and has outperformed transformer-based methods such as CSDI (Tashiro et al., 2021). The architecture consists of 4 hidden layers, four residual layers, and uses residual, skip, and hidden dimensions of size 64. The model employs embeddings for diffusion steps, with dimensions of 32 (input), 64 (middle), and 64 (output). S4 layers are configured with a maximum sequence length of 100, a state dimension of 64, and are used bidirectionally with layer normalisation and no dropout.
@@
To ensure fairness, we re-implement all diffusion baselines: SSSD (Alcaraz and Strodthoff, 2023), TimeWeaver (Narasimhan et al., 2024), TSDiff (Kollovieh et al., 2024), and TimeAutoDiff (Suh et al., 2024), using the same backbone architecture described above. SSSD is trained to impute randomly masked signal rows conditioned on the observed rows and metadata. We uniformly sample a number of timesteps between 0 and the total sequence length to randomly mask out during training. The model then trains to reconstruct the masked signals using the observed rows and the metadata. TimeWeaver, by design, conditions solely on metadata. To match this, we train it by masking out all signal rows, so that it learns to generate signals from just the metadata. We extend TSDiff to condition on metadata using the same metadata-conditioned denoiser as TimeWeaver. The observed signals are then conditioned on directly at inference time using self-guidance (see Sec. 2.2). We tried different scales for tuning TSDiff’s guidance strength: 0.0, 0.5, 1.0, and 2.0, and report results for 0.5, since it had the best overall performance (see Appendix B). For fairness, we adapt TimeAutoDiff from its original latent diffusion architecture to operate in real space using our shared backbone. It uses classifier-free guidance that conditions on labels (see Sec. 2.2), interpolating using two model variants. We adapt it to conditions on both the metadata and signal observations. One branch uses only metadata while the other conditions on both metadata and signal observations, like the SSSD baseline. We train each variant separately, and at inference, we interpolate between them using a guidance strength of 3.0, following prior work (Ho and Salimans, 2021; Suh et al., 2024).
The comparison also includes the GAN-based method TimeGAN (Yoon et al., 2019), which comprises five networks: encoder, recovery (decoder), generator, supervisor, and discriminator. We extend the base model to incorporate metadata. Each network uses a three-layer GRU (Cho, 2014) with fully connected output heads.
For the diffusion backbones, a linear noise schedule is adopted, with 200 noising and denoising steps, using $\beta_{0}=0.0001$ up to $\beta_{T}=0.02$. The guidance strength $\eta$ is fixed at 0.1 across all datasets and task configurations of WaveStitch. Unless otherwise specified, we report WaveStitch’s results using a stride of 8, since it had the best performance overall.The diffusion backbones are trained using the MSE loss between the estimated and actual noise whereas TimeGAN is optimised using the L1-regularised reconstruction loss between generated and ground truth signals. All models are trained for 300 epochs using a mini-batch size of 1024 and a learning rate of $10^{-4}$. The datasets are divided using sliding windows of 32 timesteps (stride 1) for training, and tunable stride lengths at inference.
@@ 5.3. Comparison with SOTA Generators (RQ1)
Table 4 shows the average MSE between generated and ground truth signal values over five trials. We use multiple metrics to assess performance—MSE for imputation accuracy, and quality metrics ACD and $x$-Corr for generative tasks.
Table 6. Autoregressive (WaveStitchAR) versus parallel generation across various strides (1, 8, 16, 32), datasets and tasks.
@@
Table 6 shows the MSE, average runtime in seconds and the total number of denoiser calls (hardware-agnostic metric) of the parallel (WaveStitch) and autoregressive (WaveStitchAR) implementations under varying stride lengths. We average all results over five trials, using a window size of 32 consistent across all variants, with 200 denoising steps and a mini-batch size of 1024. We also report TimeGAN results to contextualise the tradeoff between generation quality and speed against a fast generative baseline.
Overall, the MSE scores between the parallel (WaveStitch) and autoregressive (WaveStitchAR) variants are comparable, with the parallel version generally doing better. This result indicates that parallel synthesis with stitching maintains similar coherence to sequential generation. While both versions share the same model and sliding window setup, they differ in how they apply conditioning over time. In WaveStitchAR, each generated window is used as context for the next one, so early errors can propagate unchecked, degrading quality for the subsequent windows. In contrast, WaveStitch imposes conditions ”softly” via the conditional loss, iteratively refining all windows together during denoising. This joint optimisation improves robustness to early mistakes, giving WaveStitch a slight edge over WaveStitchAR.
We also observe that the parallel approach has a much faster runtime. For example, in task R in the PE dataset, WaveStitch-16 (stride 16) achieves an average speedup of $\mathbf{166.48}$ (i.e., $\frac{1566.418}{9.409}$) over its sequential counterpart WaveStitchAR-16. It also achieves a lower MSE and reduces the function calls from 160200 to just 200. The runtime generally decreases with larger strides, resulting in fewer windows, so it has fewer denoiser calls. However, for shorter sequences like AT, where nearly all windows fit within one mini-batch (1024), the stride length has little impact on runtime.
@@
The speedup does not reach the theoretical limit of the mini-batch size $b$ in practice (see Sec. 2). This reduction is possibly due to the under-utilisation of hardware and other overheads. For example, consider the MT dataset, with a sequence length of 7949 timesteps (see Table 2). With a stride of 8 and a window size of 32, the number of windows is roughly $(7949-32)/8\approx 990$. As this is less than the mini-batch size of 1024, the autoregressive model makes roughly [^4] The exact total is 198600 due to a minor implementation detail where additional timesteps are added for conditional context. $990\times 200=198000$ denoiser calls. This is lower than the worst-case with $1024\times 200=204800$ calls, so we have a lower speedup.
When analysing the speed–quality tradeoff, diffusion models consistently have better quality (lower MSE) than TimeGAN despite slower speeds. This difference stems from the core design: diffusion iteratively denoises samples, correcting errors at each step, while GANs use a single step, which is unstable.
@@ 5.5. Ablation: Effect of Stitching (RQ3)
Table 7. Comparing conditioning methods on imputation accuracy (MSE)
@@
Table 7 evaluates the imputation accuracy (MSE) using five trials, of three conditioning strategies: (1) WaveStitch with RePaint-based conditioning (Lugmayr et al., 2022) and enforced overlap-coherence (Appendix A); (2) WaveStitch with only self-guidance on observations without stitching; and (3) WaveStitch with self-guidance and stitching.
Results show that using stitching achieves the best performance in nearly all cases, while relying only on self-guidance performs the worst. RePaint is competitive, but generally loses to ours. Overall, the stitch loss improves performance by up to 35.43% compared to just self-guidance (AT, I task) and up to 28.36% over RePaint (RS, B task). This result aligns with Proposition 3.1 (Sec. 3.5), which explains how using self-guidance and stitching increases the likelihood of generating realistic samples.
@@ 5.7. Random Imputation Tasks (RQ5)
Table 9. MSE for varying misingness ratios
@@
By default, we evaluate WaveStitch on three imputation granularities: R (coarse-grained), I (intermediate), and B (fine-grained), as detailed in Sec. 5.1. However, this does not limit its generalizability to other scenarios. To show this, we also compare the accuracies (MSE) on random imputation scenarios. These are constructed by randomly masking out a certain percentage of signal rows (across signal features) while only keeping metadata intact. Results are shown in Table 9. As expected, increasing the proportion of missing information generally degrades performance. On comparing these results to the R tasks (root level) in Table 4, we find that MSE scores remain consistently better than the R tasks, which have 100% missing signals (hardest scenario). This highlights that WaveStitch is not limited to specific missingness configurations and is generalisable to other scenarios.
@@
5.8. Visual Assessment of Generated Data
We see that methods conditioning on observed signals (TSDiff and WaveStitch) better resemble ground truth signals under finer-grained tasks. In contrast, TimeWeaver does significantly worse on bottom-level tasks since it does not condition on the observed signals. Cross-feature correlations (see 7(b)) further highlight this trend, with WaveStitch and TSDiff maintaining better inter-feature dependencies for fine-grained tasks. These highlight WaveStitch’s strength in capturing key temporal and structural patterns.
6. Related Work
Table 10 summarises the related work from time series generation methods and adjacent domains. Time series generative models vary in their backbone architecture (GANs versus diffusion); their ability to condition on both metadata and observations; and the conditioning strategy (training versus inference-time).
Time Series Models: Generative models have evolved in recent years. TimeGAN (Yoon et al., 2019) pioneered GAN-based architectures for time series generation, but focuses on unconditional synthesis without leveraging metadata or the observed signals. In contrast, more recent methods employ conditioning to utilise additional information for guiding the generative process. For instance, TimeWeaver (Narasimhan et al., 2024) trains a conditional diffusion backbone by generating the entire signal just from the metadata. However, it ignores the observed signals, despite their criticality for fine-grained imputation tasks. On the other hand, TimeGrad (Rasul et al., 2021) ignores metadata and only handles forecasting tasks by conditioning only on signals from the preceding timesteps, making it unsuitable for imputation at arbitrary timesteps. Moreover, it maintains temporal coherence via autoregression, which is slow due to sequential generation. SSSD (Alcaraz and Strodthoff, 2023) includes the conditional masks as inputs during training. While this helps impute values at arbitrary positions, there is a risk of overfitting on the masking patterns seen during training. TSDiff (Kollovieh et al., 2024) removes this risk by conditioning on the observed signal patterns directly at inference through self-guidance. However, self-guidance only conditions on the observations within a given time window, ignoring dependencies across overlaps. TimeAutoDiff (Suh et al., 2024) sidesteps this cross-window dependency problem by generating the entire output in one go using a latent transformer-based diffusion model (Rombach et al., 2022) and classifier-free guidance (Ho and Salimans, 2021) for conditioning on labels. However, this design restricts it to fixed-length sequences and is impractical: longer sequences significantly increase the computational load on the self-attention layers, hurting scalability (Keles et al., 2023). Moreover, adapting classifier-free guidance from conditioning on discrete labels to conditioning on the continuous and arbitrarily observed signals is challenging.
@@ 6. Related Work
In summary, many existing related works rely on minimal or rigid conditioning, ignore time dependencies, or are inefficient. Instead, WaveStitch offers a lightweight, inference-time solution that flexibly accommodates diverse conditioning signals and maintains temporal consistency without retraining or sacrificing parallelism.

### v3

@@ 5. Evaluation
RQ3: Does stitching improve performance compared to using only self-guidance or RePaint (Lugmayr et al., 2022; Chen et al., 2023)?
RQ4: What is the impact of alternative conditional loss formulations on the generation quality? RQ5: Does WaveStitch generalise to random imputation tasks with varying missingness ratios?
System Configuration. We use a 12-core AMD Ryzen 9 5900 processor with an NVIDIA RTX 3090 GPU. For software, we use CUDA 12.1, Ubuntu 20.04.6, and Pytorch 2.2.1 with Python 3.12.
@@ 5.1. Datasets and Evaluation Framework
Datasets. We use public multivariate and univariate datasets from diverse domains: BeijingAirQuality (BQ) (Chen, 2017), MetroTraffic Volume (MT) (Hogue, 2019), PanamaEnergy (PE) (Aguilar Madrid, 2021), RossmanSales (RS) (FlorianKnauer and Cukierski, 2015), and AustraliaTourism (AT) (Blanche, 2020). Each dataset includes metadata, grouped in the order specified in Table 2, reflecting how the data is organised. The time series channels for each dataset are as follows. For BQ, we use all gas sensor readings, temperature, pressure, dewpoint, rainfall, and wind speed from the first six stations, sorted alphabetically; for MT, we use temperature, rainfall, snowfall, traffic volume, and cloud cover levels; for PE, we use energy data such as T2M, QV2M, TQL, and W2M; for RS, we use sales and customer totals from stores 1-10; and for AT we use the number of trips.
Evaluation Tasks. Designing evaluation scenarios for our problem is not trivial. Random or standard 80-20 splits would disrupt temporal dependencies. To address this, we place all samples with a shared root-level value (e.g., a specific year) in the test set, ensuring unseen metadata combinations at inference. This approach supports coarse-grained imputation tasks, as well as fine-grained or mid-level tasks, by specifying lower-level metadata features within the test set.
To evaluate the model’s performance across different scenarios, we use a structured approach that conditions metadata at three levels: Root (R), Intermediate (I), and Bottom (B), each corresponding to a different level in the grouping hierarchy and reflecting varying degrees of missing signal information and task granularity (see Table 2). Root Level (R) conditions fix the root-level metadata feature (e.g. Year in a Year-¿Month-¿Day hierarchy), creating tasks missing contiguous signal blocks (e.g., 4(a)). Intermediate Level (I) tasks restrict mid-level metadata features, giving missing blocks of intermediate size (e.g., 4(b)). Finally, Bottom Level (B) tasks restrict the last metadata feature, yielding fine-grained imputation scenarios (e.g., 4(c)). As shown in Table 2, both I and B additionally condition on the root-level feature to ensure that samples remain within the test set. While these configurations represent only a subset of possible scenarios, they cover a broad range of missingness patterns, allowing us to evaluate the model under various conditions. For example, this setup can also capture random missingness by combining multiple (B) conditions.
@@ 5.2. Model Implementation and Baselines
We extend the implementations of several baselines from Table 10 for conditional generation. For all diffusion models, including ours, we use SSSD (Alcaraz and Strodthoff, 2023) as the denoiser backbone for consistency, which utilises structured state space layers (S4 layers (Smith et al., 2023)) and has outperformed transformer-based methods such as CSDI (Tashiro et al., 2021). The architecture consists of 4 hidden layers, four residual layers, and uses residual, skip, and hidden dimensions of size 64. The model employs embeddings for diffusion steps, with dimensions of 32 (input), 64 (middle), and 64 (output). S4 layers are configured with a maximum sequence length of 100, a state dimension of 64, and are used bidirectionally with layer normalisation and no dropout.
To ensure fairness, we re-implement all diffusion baselines: SSSD (Alcaraz and Strodthoff, 2023), TimeWeaver (Narasimhan et al., 2024), TSDiff (Kollovieh et al., 2024), and TimeAutoDiff (Suh et al., 2024), using the same backbone architecture described above. SSSD is trained to impute randomly masked signal rows conditioned on the observed rows and metadata. We uniformly sample a number of timesteps between 0 and the total sequence length to randomly mask out during training. The model then trains to reconstruct the masked signals using the observed rows and the metadata. TimeWeaver, by design, conditions solely on metadata. To match this, we train it by masking out all signal rows, so that it learns to generate signals from just the metadata. We extend TSDiff to condition on metadata using the same metadata-conditioned denoiser as TimeWeaver. The observed signals are then conditioned on directly at inference time using self-guidance (see Sec. 2.2). We tried different scales for tuning TSDiff’s guidance strength: 0.0, 0.5, 1.0, and 2.0, and report results for 0.5, since it had the best overall performance. For fairness, we adapt TimeAutoDiff from its original latent diffusion architecture to operate in real space using our shared backbone. It uses classifier-free guidance that conditions on labels (see Sec. 2.2), interpolating using two model variants. We adapt it to conditions on both the metadata and signal observations. One branch uses only metadata while the other conditions on both metadata and signal observations, like the SSSD baseline. We train each variant separately, and at inference, we interpolate between them using a guidance strength of 3.0, following prior work (Ho and Salimans, 2021; Suh et al., 2024).
We also compare against the GAN-based method TimeGAN (Yoon et al., 2019), which consists of five networks: encoder, recovery (decoder), generator, supervisor, and discriminator. We extend the base model to incorporate metadata. Each network uses a three-layer GRU (Cho et al., 2014) with fully connected output heads.
For the diffusion backbones, we use a linear noise schedule with 200 noising and denoising steps, using $\alpha_{1}=0.9999$ and $\alpha_{T}=0.98$. The guidance strength $\eta$ is fixed at 0.1 for all datasets and task configurations. By default, we report WaveStitch’s results using a stride of 8, as it achieved the best overall performance. We trained the diffusion backbones using the MSE loss between the estimated and actual noise, and we optimised TimeGAN using the L1-regularised reconstruction loss between the generated and ground-truth signals. We train all models for 300 epochs using a mini-batch size of 1024 and a learning rate of $10^{-4}$. For consistency, we divided all the datasets using sliding windows of 32 timesteps (stride 1) for training, and tunable stride lengths at inference.
@@ 5.3. Comparison with SOTA Generators (RQ1)
Table 3 shows the average MSE between the generated and ground truth signals over five trials. We use multiple metrics to assess performance: MSE for imputation accuracy, and quality metrics ACD and $x$-Corr for generative tasks.
Table 6. Autoregressive (WaveStitchAR) versus parallel generation across various strides (1, 8, 16, 32), datasets and tasks. The best scores are marked in bold.
@@
Table 6 shows the MSE, average runtime (seconds) and the total number of denoiser calls of the parallel (WaveStitch) and autoregressive (WaveStitchAR) implementations under varying stride lengths. We average all results over five trials, using a window size of 32 consistent across all variants, with 200 denoising steps and a mini-batch size of 1024. We also report TimeGAN results to contextualise the tradeoff between generation quality and speed against a fast generative baseline.
Overall, the MSE scores between the parallel (WaveStitch) and autoregressive (WaveStitchAR) variants are comparable, with the parallel version generally doing better. This result indicates that parallel synthesis with stitching maintains similar coherence to sequential generation. While both versions share the same model and sliding window setup, they differ in how they apply conditioning over time. In WaveStitchAR, each generated window is used as context for the next one, so early errors can propagate unchecked, degrading quality for the subsequent windows. In contrast, WaveStitch imposes conditions “softly” via the conditional loss, iteratively refining all windows together during denoising. This joint optimisation improves robustness to early mistakes, giving WaveStitch a slight edge over WaveStitchAR.
The parallel approach also has a much faster runtime. For example, in task R in the PE dataset, WaveStitch-16 (stride 16) achieves an average speedup of $\mathbf{166.48}$ (i.e., $\frac{1566.418}{9.409}$) over its sequential counterpart WaveStitchAR-16. It also achieves a lower MSE and reduces the function calls from 160200 to just 200. The runtime generally decreases with larger strides, resulting in fewer windows, thus resulting in fewer denoiser calls. However, for shorter sequences like AT, where nearly all windows fit within one mini-batch (1024), the stride length has little impact on runtime.
@@
The speedup does not reach the theoretical limit of the mini-batch size $b$ in practice (see Sec. 6), possibly due to the under-utilisation of hardware and other overheads. For example, consider the MT dataset, with a sequence length of 7949 timesteps (see Table 2). For window size 32 and stride 8, the number of windows is roughly $(7949-32)/8\approx 990$. As this is less than the mini-batch size of 1024, the autoregressive model makes roughly [^4] The exact total is 198600 due to a minor implementation detail where additional timesteps are added for conditional context. $990\times 200=198000$ denoiser calls. This is lower than the worst-case with $1024\times 200=204800$ calls, so we have a lower speedup.
When analysing the speed–quality tradeoff, diffusion models consistently have lower MSEs than TimeGAN despite slower speeds. This difference stems from the core design: while diffusion iteratively corrects errors at each step, GANs generate using a single step, which is unstable.
@@ 5.5. Ablation: Effect of Stitching (RQ3)
Table 7 evaluates the imputation accuracy (MSE) using five trials, of three conditioning strategies: (1) WaveStitch with RePaint-based conditioning (Lugmayr et al., 2022) and enforced overlap-coherence; (2) WaveStitch with only self-guidance on observations without stitching; and (3) WaveStitch with self-guidance and stitching. Results show that using stitching achieves the best performance in nearly all cases, while relying only on self-guidance performs the worst. RePaint is competitive, but generally loses to ours. Overall, the stitch loss improves performance by up to 35.43% compared to just self-guidance (AT, I task) and up to 28.36% over RePaint (RS, B task). This result directly aligns with Proposition 3.1 (Sec. 3.5), which explains how using self-guidance and stitching increases the likelihood of generating realistic samples.
Table 7. Comparing conditioning methods on imputation accuracy (MSE).
@@ 5.7. Random Imputation Tasks (RQ5)
By default, we evaluate WaveStitch on three imputation granularities: R (coarse-grained), I (intermediate), and B (fine-grained), as detailed in Sec. 5.1. However, this does not limit its generalisability to other scenarios. To show this, we also compare the accuracies (MSE) on random imputation scenarios. These tasks are constructed by randomly masking out a certain percentage of signal rows while keeping the metadata intact. Results are shown in Table 9. As expected, increasing the proportion of missing information generally degrades performance. On comparing these results to the R tasks (root level) in Table 3, we find that MSE scores remain consistently better than the R tasks, which have 100% missing signals (hardest scenario). This highlights that WaveStitch is not limited to specific missingness configurations and is generalisable to other scenarios.
Table 9. MSE for varying missingness ratios
@@
5.8. Visual Assessment of Generated Data
We see that methods conditioning on observed signals (TSDiff and WaveStitch) better resemble ground truth signals under finer-grained tasks. In contrast, TimeWeaver does significantly worse on bottom-level tasks since it does not condition on the observed signals. Cross-feature correlations (see 7(b)) further highlight this trend, with WaveStitch and TSDiff maintaining better inter-feature dependencies for fine-grained tasks. These results highlight WaveStitch’s strength in capturing key temporal and structural patterns.
6. Related Work
Table 10 summarises the related works from time series generation and adjacent domains. These works vary in their backbone architectures (GANs versus diffusion); their ability to condition on both metadata and observations; and their conditioning strategies (training versus inference-time).
Time Series Models: Generative models have evolved in recent years. TimeGAN (Yoon et al., 2019) pioneered GAN-based architectures for time series generation, but focuses on unconditional synthesis without leveraging metadata or the observed signals. In contrast, more recent methods employ conditioning to utilise additional information for guiding the generative process.
TimeWeaver (Narasimhan et al., 2024) trains a conditional diffusion backbone by generating the entire signal just from the metadata. However, it ignores the observed signals, despite their criticality for fine-grained imputation tasks. In contrast, TimeGrad (Rasul et al., 2021) ignores metadata but only handles forecasting tasks by conditioning on historical signals from preceding timesteps, which is unsuitable for imputation at arbitrary timesteps. Moreover, it maintains temporal coherence via autoregression, which is slow due to sequential generation. SSSD (Alcaraz and Strodthoff, 2023) includes the conditional masks as inputs during training. While this enables imputing values at arbitrary timesteps, it risks overfitting on the masking patterns seen during training. TSDiff (Kollovieh et al., 2024) removes this risk by conditioning on the observed signal patterns directly at inference through self-guidance. However, self-guidance only conditions on the observations within a given time window, ignoring dependencies across overlaps. TimeAutoDiff (Suh et al., 2024) sidesteps this cross-window dependency problem by generating the entire output in one go using a latent transformer-based diffusion model (Rombach et al., 2022) and classifier-free guidance (Ho and Salimans, 2021) for conditioning on labels. However, this design restricts it to fixed-length sequences and is impractical: longer sequences significantly increase the computational load on the self-attention layers, hurting scalability (Keles et al., 2023). Moreover, adapting classifier-free guidance from conditioning on discrete labels to conditioning on the continuous and arbitrarily observed signals is challenging.
@@ 6. Related Work
In summary, many existing related works rely on minimal or rigid conditioning, ignore temporal dependencies, or are inefficient. Instead, WaveStitch offers a lightweight, inference-time solution that flexibly accommodates diverse conditioning signals and maintains temporal consistency without retraining or sacrificing parallelism.